OpenStudy (anonymous):

Graph y= -2x ^{2} +6 Describe what you see. I don't know how to graph that?

OpenStudy (anonymous):

how old r u

OpenStudy (anonymous):

lol does that really matter? I just need help with the problem.

OpenStudy (anonymous):

@ASVPJOHNNY

OpenStudy (anonymous):

This is an equation of the standard form for a parabola: y=A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex. For given equation: vertex:(0,6) negative lead coefficient means parabola opens downwards y-intercept set x=0 y=6 .. x-intercepts: set y=0 2x^2=6 x^2=3 x=±√3 .. you now have the vertex(0,6), y-intercept(0,6) and two x-intercepts(±√3,0) with which you can easily draw the graph of given function. --------------------------------------------------------------------------------